OpenStudy (anonymous):

Find an equation in the format of f(x)=b^x given the point (2,e)

OpenStudy (anonymous):

@ (2,e): e=b^2 then: ln (e)=ln (b^2) e=2 ln(b) ln(b)=1/2 e \[b=e ^{1/2}\]

OpenStudy (anonymous):

now you know b, then: \[f(x)=(e ^{1/2})^{x}=e ^{x/2}\]

OpenStudy (anonymous):

I don't think that's right.

OpenStudy (anonymous):

why not...?

OpenStudy (anonymous):

We haven't used natural logs yet. That's the next section.

OpenStudy (anonymous):

.... the answer is right - put the value of the given point (2,e)=> you'll have a true statement.

OpenStudy (anonymous):

\[e = b^2 \implies \sqrt{e} = b\]

OpenStudy (anonymous):

\[\implies f(x) = (\sqrt{e})^x\]

OpenStudy (anonymous):

i thing you can use: e=b^2 then taking squire root both sides: b=e^(1/2)

OpenStudy (anonymous):

Yes, that's right, I answered this already in another post.